The student council at your school consists of 22 members. In how many ways can you select a committee of five members if
a) the president and the secretary must be included?
b) the vice-president must be excluded?
1
Expert's answer
2020-03-06T16:04:00-0500

There are 22 members. A commitee of 5 members is needed.

a) the president and the secretary included. 20 members remain and the committee will need 3 members to be selected.

C(n, r) ="\frac {n!} {(n-r)! r!}"

C(20,3)= "\frac {20!}{(20-3)!3!}"

=1140 ways

b) vice president must be excluded

Eligible members will be 21 and 5 members need to be selected to form a commitee

C(n, r) = "\frac {n!} {(n-r)! r!}"

C(21,5)="\frac {21!}{(21-5)!5!}"

=20349 ways